OpenStudy (beth12345):

The gravitational field strength is also governed by an inverse square law. What is the gravitational field strength 200km above the Earth's surface, at the altitude of many manned space flights?

OpenStudy (beth12345):

The answer I'm given is 9.2N/kg but U don't know how to get it. I also have an equation sheet with the following equations...\[Fg=m \times g\]\[F=G \times( (m \ _{1} \times m _{2})/r ^{2})\]\[F _{fr}=\mu \times F _{N}\]\[F _{net}=m \times a\]\[p=m \times v\]\[\Delta p= F \times \Delta t\]\[F = k \times x\] but I also don't know what one to use, if someone could help me, that would be nice :)

OpenStudy (nottim):

I think you should just look for what's given. Once you get that, see if only one variable is left. If it is, then you got your equation.

OpenStudy (nottim):

ok?

OpenStudy (anonymous):

i think u should go for equation 1 2 of force as in question it is talking about the gravitational field.